The kinetic energy of a charged particle decreases by 10 J as it moves from a point at potential 100 V to a point at potential 200 V. Find the charge on the particle.

Grade:11

1 Answers

Kevin Nash
9 years ago
Sol. K.C. decreases by 10 J. Potential = 100 v to 200 v. So, change in K.E = amount of work done ⇒ 10J = (200 – 100) v × q base 0 ⇒ 100 q base 0 = 10 v ⇒ q base 0 = 10/100 = 0.1 C